### v1

# Comparing Left and Right Quotient Sets in Groups

Abstract.
For a finite subset $A$ of a group $G$, we define the right quotient set and the left quotient set of $A$, respectively, as
$AA^{-1}$ $~\coloneqq~\{a_{1}a_{2}^{-1}:a_{1},a_{2}\in A\},$
$A^{-1}A$ $~\coloneqq~\{a_{1}^{-1}a_{2}:a_{1},a_{2}\in A\}.$
While the right and left quotient sets are equal if $G$ is abelian, subtleties arise when $G$ is a nonabelian group, where the cardinality difference $|AA^{-1}|-|A^{-1}A|$ may be take on arbitrarily large values. Using the results of Martin and O’Bryant on the cardinality differences of sum sets and difference sets in $\mathbb{Z}$, we prove in the infinite dihedral group, $D_{\infty}\cong\mathbb{Z}\rtimes\mathbb{Z}/2\mathbb{Z}$, every integer difference is achievable. Further, we prove that in $F_{2}$, the free group on $2$ generators, an integer difference is achievable if and only if that integer is even, and we explicitly construct subsets of $F_{2}$ that achieve every even integer. We further determine the minimum cardinality of $A\subset G$ so that the difference between the cardinalities of the left and right quotient sets is nonzero, depending on the existence of order $2$ elements in $G$. To prove these results, we construct difference graphs $D_{A}$ and $D_{A^{-1}}$ which encode equality, respectively, in the right and left quotient sets. We observe a bijection from edges in $D_{A}$ to edges in $D_{A^{-1}}$ and count connected components in order to obtain our results on cardinality differences $|AA^{-1}|-|A^{-1}A|$.
This research was supported by the National Science Foundation, under NSF Grant DMS2241623, Amherst College, Princeton University, the University of Wisconsin, the University of Michigan, Williams College, and the Finnerty Fund. We thank the participants of the 2025 Integers Conference for conversations indirectly inspiring this work.
1. Introduction
Note: This paper is dedicated with thanks to Carl Pomerance and Mel Nathanson. The genesis for this work came from conversations the fifth named author had at the Integers Conference in Georgia in 2025 on Results in Additive & Elementary Number Theory Inspired by Carl and Mel. His presentation described how much of his mentoring of students has been influenced by each, in particular problems in the orbit of MSTD sets (from Mel) and multiplicative and quotient structures (from Carl). This led to springboard problems for the SMALL 2025 REU, which led to the work below.
1.1. Background
Given a subset $A$ of $N]~\coloneq~\{1,...,N\}$, the sumset and difference sets are defined as
$A+A$ $~\coloneqq~\{a_{1}+a_{2}:a_{1},a_{2}\in A\},$
$A-A$ $~\coloneqq~\{a_{1}-a_{2}:a_{1},a_{2}\in A\}.$
A natural comparison arises between the cardinalities of the sum and difference sets. Our set $A$ is said to be sum dominated or MSTD (more sums than differences) if |A+A|>|AA||A+A|>|A-A| and difference dominated (MDTS) if |AA|>|A+A||A-A|>|A+A|. One might expect that the difference set would have a greater cardinality as addition is commutative in $\mathbb{Z}$ while subtraction is not. In particular, if we consider a pair of distinct elements $a_{1},a_{2}\in A$, $a_{1}-a_{2}$ and $a_{2}-a_{1}$ are distinct elements in the difference set, while $a_{1}+a_{2}=a_{2}+a_{1}$ is a single element in the sumset. However, it is possible to have a set that is sum dominated. The earliest example of a MSTD set was discovered by Conway in 1960s: $\{0,2,3,4,7,11,12,14\}$. Surprisingly, in contrast to the intuition that MSTD sets should form a vanishing proportion of subsets of $[N]$ as $N$ grows large, in 2006 Martin and O’Bryant [MO06] proved that the proportion of subsets of $N]\subseteq\mathbb{Z}$ which are MSTD does not vanish as $N\to\infty$. Since then, extensive research has expanded the classical MSTD problem to various settings, including higher dimensions and various families [KM22, CLMS20, CLMS19].
One natural extension of the study of MSTD sets amongst subsets of $\mathbb{Z}$ is to ask this question for groups in general. We primarily focus on the free group $F_{2}$ on two generators and the infinite dihedral group $D_{\infty}\cong\mathbb{Z}\rtimes\mathbb{Z}/2\mathbb{Z}$.
Definition 1.1.
Let $A$ be a set of generators with $|A|=n$. The free group on $A$, denoted $F(A)$ or $F_{n}$, is the group consisting of all reduced words over the alphabet $A\cup A^{-1}=\left\{x:x\in A\right\}\cup\left\{x^{-1}:x\in A\right\}$ under the operation of concatenation followed by reduction, where reduction means canceling adjacent inverse pairs.
Following previous work by [ACJ+22, MV14, Nat07, Zha10], we look at sumsets and difference sets in general groups. Let $A\subset G$ be a finite subset of a group $G$ equipped with the operation $\times:(x,y)\mapsto xy$ and the inverse map $^{-1}:x\mapsto x^{-1}$. Although some authors use the terminology sumset and difference set in this context [ACJ+22, MV14, Nat07, Zha10], we will refer to the product set and the quotient set of $A$, which are defined, respectively as
$AA$ $~\coloneqq~\{xy:x,y\in A\},$
$AA^{-1}$ $~\coloneqq~\{xy^{-1}:x,y\in A\}.$
More generally, let $G$ be any group. Let $A=\left\{a_{1},\dots,a_{n}\right\}\subseteq G$. Define $A^{-1}\coloneqq\left\{a_{1}^{-1},\dots,a_{n}^{-1}\right\}$. We consider the right and left quotient sets of $A$, defined, respectively, by
$AA^{-1}~\coloneqq~\left\{a_{i}\cdot a_{j}^{-1}:a_{i},a_{j}\in A\right\},$
$A^{-1}A~\coloneqq~\left\{a_{i}^{-1}\cdot a_{j}:a_{i},a_{j}\in A\right\}.$
In the classical MSTD problem, one natural area of study has been the cardinality of the smallest MSTD set. Hegarty showed that the smallest such example in $\mathbb{Z}$ has cardinality $8$, and is unique up to affine translation [Heg07]. Another natural question asks for the set of values |A+A||AA||A+A|-|A-A| can attain across all $A\subset[N]$. Martin and O’Bryant proved, over finite subsets of $\mathbb{Z}$, that this difference achieves all integers [MO06].
Inspired by this, we consider similar questions about left and right quotient sets. First note that the problem of constructing a set $A$ where $|AA^{-1}|>|A^{-1}A|$ is the same as the problem of $|AA^{-1}|<|A^{-1}A|$ (by replacing $A$ with $A^{-1}$). Therefore, a more natural question to ask is the smallest $A$ where the left and right quotient sets are not equal. Moreover, inspired by [MO06], we give results on what values $|AA^{-1}|-|A^{-1}A|$ can take as $A$ ranges over finite subsets of a group $G$.
Example $1.$
While there has been some work done generalizing in a abelian group setting [MV14, Nat07, Zha11], our questions are only relevant for non-abelian groups. Indeed, if $G$ is abelian, then all $A\subseteq G$ have $AA^{-1}=A^{-1}A$, so the only possible value of $|AA^{-1}|-|A^{-1}A|$ is 0.
1.2. Notation and Main Results
Theorem 1.1.
Let $G$ be a group with no elements of order 22.111While we use the language of “a group with no elements of order 22”, we remark that Tao refers to such a group as a “a $2$-torsion-free group” in [Tao08]. Let $A\subseteq G$ be a finite subset. Then $|AA^{-1}|-|A^{-1}A|$ is even.
This result does not hold in the context of groups with elements of order $2$. In fact, we can construct examples where odd values of $|AA^{-1}|-|A^{-1}A|$ are achieved.
Example 1.2.
Let $G=D_{\infty}\coloneqq\left\langle r,s\mid sr=r^{-1}s,s^{2}=e\right\rangle$ be the infinite dihedral group. For every $n\in\mathbb{Z}$, there exists a subset $A_{n}\subseteq D_{\infty}$ such that $|A_{n}A_{n}^{-1}|-|A_{n}^{-1}A_{n}|=n$. Indeed, let $B\subseteq\mathbb{Z}$ be a finite subset to be determined later. Consider the subset
$A~\coloneqq~\{r^{b}:b\in B\}~\cup~\{sr^{b}:b\in B\}~\subseteq~D_{\infty}.$
We see that
$AA^{-1}~=$~ $\underbrace{\left\{r^{b-b^{\prime}}:b,b^{\prime}\in B\right\}}_{A_{1}}~\sqcup~\underbrace{\left\{sr^{b^{\prime}-b}:b,b^{\prime}\in B\right\}}_{A_{2}},$
$A^{-1}A~=$~ $\underbrace{\left\{r^{b^{\prime}-b}:b,b^{\prime}\in B\right\}}_{A_{1}^{\prime}}~\sqcup~\underbrace{\left\{sr^{b^{\prime}+b}:b,b^{\prime}\in B\right\}}_{A_{2}^{\prime}}.$
Notice that $A_{1}=A_{1}^{\prime}$, and $|A_{2}|=|B-B|$ and $|A_{2}^{\prime}|=|B+B|$. Therefore,
$|AA^{-1}|-|A^{-1}A|~=~|B-B|-|B+B|.$
By [MO06, Theorem 4], the latter difference ranges over every integer.
Considering in particular the free group $G=F_{2}$, we can construct examples where every even integer is achieved.
Theorem 1.2 ($F_{2}$ achieves all even possible differences).
For all $n\in\mathbb{Z}$, there exists a set $A_{n}\subseteq F_{2}$ such that $|A_{n}A_{n}^{-1}|-|A_{n}^{-1}A_{n}|~=~2n$.
The construction for the previous theorem uses a subset $A\subseteq F_{2}$ of cardinality $5$ that satisfies $|AA^{-1}|\neq|A^{-1}A|$. The following theorem shows this construction is optimal with respect to the size of $A$ for a more general class of groups.
Theorem 1.3.
Let $G$ be a group. Let $A\subseteq G$ be a finite subset and suppose that $|AA^{-1}|\neq|A^{-1}A|$. Then
  • without any further assumptions, $|A|\geq 4$, and
  • if $G$ is a group with no elements of order $2$, then $|A|\geq 5$.
A brute force search on groups of small order show that the bound $|A|\geq 4$ is sharp (see 2.2). The main tool we use to prove this result is a graph associated to the left (and right) quotient sets, which we call the difference graph. The cardinalities of $AA^{-1}$ and $A^{-1}A$ can be interpreted as the number of connected components on these graphs. Making use a bijection of edges between these graphs, we perform an argument based on the properties of this graph to prove that when $|A|\leq 3$ (resp. $|A|\leq 4$ when $A$ has no elements of order $2$), the number of connected components does not change under the bijection of edges.
2. Left vs. Right Quotient Sets
2.1. Graph Construction
To prove our results, we define the difference graph $D_{A}$ of a finite subset $A\coloneqq\{a_{1},a_{2},...,a_{n}\}\subseteq G$. The graph $D_{A}=(V,E)$ is defined as follows.
  1. (1)
    The vertex set is given by $V\coloneqq[n]\times[n]$.
  2. (2)
    The edge set $E(D_{A})$ is given by the relation
    $(i,j)\sim(k,\ell)\iff a_{i}a_{j}^{-1}=a_{k}a_{\ell}^{-1}.$
The difference graph is directed and not simple (we allow self-loops).
Similarly, for $D_{A^{-1}}$, we have the edge relation
$(i,j)\sim(k,\ell)\iff a_{i}^{-1}a_{j}=a_{k}^{-1}a_{\ell}.$
We first note the following basic facts about $D_{A}$.
Lemma 2.1 (Properties of $D_{A}$).
Let $i,j,k,\ell\in[n]$.
  1. (1)
    $(i,j),(k,\ell)]\in E(D_{A})\iff[(j,i),(\ell,k)]\in E(D_{A})$.
  2. (2)
    The following types of edges are not present in $E(D_{A})$.
    1. (a)
      $[(i,j),(k,k)]$, an edge connecting to the diagonal, provided that $i\neq j$.
    2. (b)
      $[(i,j),(i,k)]$ (or $[(j,i),(k,i)]$, but this is handled by (1)), an edge connecting vertices on the same axis.
    3. (c)
      If $G$ has no elements of order $2$, $[(i,j),(j,i)]$, an edge connecting to its symmetric pair, provided that $j\neq i$.
  3. (3)
    $(i,i),(j,j)]\in E(D_{A})$.
  4. (4)
    If $C$ is a connected component in $D_{A}$, then $C$ is a clique.
Proof.
(1) Suppose $a_{i}a_{j}^{-1}=a_{k}a_{\ell}^{-1}$. Then
$a_{j}a_{i}^{-1}~=~(a_{i}a_{j}^{-1})^{-1}~=~(a_{k}a_{\ell}^{-1})^{-1}~=~a_{\ell}a_{k}^{-1}.$
Hence, $(j,i),(\ell,k)]\in E(D_{A})$ and note the reverse follows.
(2a) Let $(i,j),(k,k)]\in E(D_{A})$ where $i\neq j$. Thus
$a_{i}a_{j}^{-1}~=~a_{k}a_{k}^{-1}~=~e\implies a_{i}~=~a_{j},$
but $i\neq j$, a contradiction.
(2b) Without loss of generality, take the edge $[(i,j),(i,k)]$. Then we have
$a_{i}a_{j}^{-1}~=~a_{i}a_{k}^{-1}\implies a_{j}~=~a_{k}.$
(2c) Let $(i,j),(j,i)]\in E(D_{A})$ where $i\neq j$.
$a_{i}a_{j}^{-1}~=~a_{j}a_{i}^{-1}~=~(a_{i}a_{j}^{-1})^{-1}\implies a_{i}a_{j}^{-1}~=~e\implies a_{i}~=~a_{j},$
contradicting our asusmption that $i\neq j$.
(3) Consider $a_{i}a_{i}^{-1}~=~e~=~a_{j}a_{j}^{-1}$, thus $(i,i),(j,j)]\in E(D_{A})$.
(4) This follows because equality is an equivalence relation. ∎
Remark 2.2.
In light of Property (4), we will refer to a connected component $C$ in $D_{A}$ consisting of $k$ elements simply by indicating its vertices $C=(a_{1},b_{1})(a_{2},b_{2})\cdots(a_{n},b_{n})$.
Remark 2.3.
Property (1) is the same as saying the “transpose” operation
${T}\colon{G}\to{G}$
$[(i,j),(k,\ell)]$ $\mapsto[(j,i),(\ell,k)]$
is a graph automorphism.
Let $\mathcal{C}(D_{A})$ be the set of connected components of $D_{A}$ and $c(D_{A})=\left\lvert\mathcal{C}(D_{A})\right\rvert$ be the number of connected components. Note that $c(D_{A})=|AA^{-1}|$ and $c(D_{A^{-1}})=|A^{-1}A|$.
Using the fact that
$a_{i}a_{j}^{-1}~=~a_{k}a_{\ell}^{-1}\iff a_{k}^{-1}a_{i}~=~a_{\ell}^{-1}a_{j},$
we obtain a bijection of edges between $D_{A}$ and $D_{A^{-1}}$ as follows:
$\begin{aligned} {\phi}\colon{E(D_{A})}&\to{E(D_{A^{-1}})}\\ [(i,j),(k,\ell)]&\mapsto[(k,i),(\ell,j)].\end{aligned}$
A priori, this map is only well-defined if we consider edges as directed (the reverse edge gets mapped to the transpose of the original edge) and allow loops (they get mapped to the diagonal). However, since $T$ is an automorphism of $D_{A}$, we may take $D_{A}$ to be undirected. This is a well-motivated map: its existence already tells us that the additive energies $\Lambda(A,A^{-1})~\coloneqq~\{\#(a_{1},a_{2},a_{3},a_{4})\in A^{4}:a_{1}a_{2}^{-1}=a_{3}a_{4}^{-1}\}$ and $\Lambda(A^{-1},A)~\coloneqq~\{\#(a_{1},a_{2},a_{3},a_{4})\in A^{4}:a_{1}^{-1}a_{2}=a_{3}^{-1}a_{4}\}$ are equal in non-commutative groups (see [Tao08]).
Example 2.1.
When $n=5$, we consider the following example.
$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$$a_{1}^{-1}$$a_{2}^{-1}$$a_{3}^{-1}$$a_{4}^{-1}$$a_{5}^{-1}$$a_{1}^{-1}$$a_{2}^{-1}$$a_{3}^{-1}$$a_{4}^{-1}$$a_{5}^{-1}$$\phi$
The former (left) graph prior to the mapping $\phi$ has $17$ connected components, and the latter (right) graph after $\phi$ is applied has $15$ connected components.
The first graph corresponds to
$a_{3}a_{1}^{-1}~=~a_{4}a_{3}^{-1}~=~a_{5}a_{2}^{-1}$
being satisfied, and no other relations between words (other than the diagonal). The set
$A~\coloneqq~\left\{x,xz,y^{-1},y^{-1}x^{-1}y^{-1},y^{-1}z\right\}\subseteq F_{3}~=~F(\left\{x,y,z\right\})$ (2.1)
satisfies this property, therefore giving us an example where $|AA^{-1}|\neq|A^{-1}A|$.
2.2. Possible Differences
The natural question that arises is: what are the possible differences between the cardinalities of the right and left quotient sets of $A\subset G$? In 1.2 we gave a construction on the infinite dihedral group $D_{\infty}$ demonstrating the difference $|AA^{-1}|-|A^{-1}A|$ achieves every possible $n\in\mathbb{Z}$. Given the restriction that there are no elements of order 2 in our group, we can conclude that $|AA^{-1}|-|A^{-1}A|$ is even. The proof of Theorem 1.1 follows.
Proof of Theorem 1.1.
It suffices to show that $c(D_{A})-c(D_{A^{-1}})$ is even. We claim that $c(D_{A})$ is odd.
We divide $\mathcal{C}(D_{A})$ into two disjoint classes.
  1. (1)
    The connected components that are fixed under $T$ (i.e., $T(C)=C$), and
  2. (2)
    the connected components which are disjoint from their image under $T$; that is, those connected components that are swapped with a distinct component under $T$.
Denote these sets, respectively, by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Note that $\left\lvert\mathcal{C}_{2}\right\rvert$ is even as each component comes in pairs.
We claim that $\mathcal{C}_{1}$ contains only the diagonal (the diagonal is a connected component by 2.1 (3, 2a)). Indeed, if any other component $C$ belongs to $\mathcal{C}_{1}$, then there exists a vertex $(i,j)$ with $i\neq j$ such that $(i,j),(j,i)]\in C$, which contradicts 2.1 (2c). Therefore, there is exactly one component of $\mathcal{C}_{1}$ while the rest of the connected components come in symmetric pairs, so $|\mathcal{C}_{1}|$ is odd. Applying the same reasoning to $c(D_{A^{-1}})$ shows it is odd, so $c(D_{A})-c(D_{A^{-1}})$ is even. ∎
We now consider the set of possible differences in $F_{2}$. In order to prove Theorem 1.2, we use the following.
Fact 2.4.
Let $m\geq 2$ be an integer. Then there exists an embedding $F_{m}\hookrightarrow F_{2}$.
Proof.
See, for example, [Löh17]. ∎
Writing out this embedding explicitly allows us to describe a subset of $F_{2}$ where $|AA^{-1}|\neq|A^{-1}A|$. Indeed, take the set $A$ given in (2.1) with the embedding $F_{3}\hookrightarrow F_{2}$ by $x\mapsto x^{2}$, $y\mapsto xy$, $z\mapsto xy^{-1}$ to get the set
$\{x^{2},x^{3}y^{-1},y^{-1}x^{-1},y^{-1}x^{-3}y^{-1}x^{-1},y^{-2}\}\subseteq F_{2}.$
Proof of Theorem 1.2.
Any set $A\subseteq\left\langle a\right\rangle\cong\mathbb{Z}$ with $a\in F_{2}$ yields the $n=0$ case. By replacing $A$ with $A^{-1}$, it suffices to prove the result for positive $n$. So we construct a family of sets $A_{n}\subseteq F_{m}$ for some $m\geq 2$ and then compose it with the embedding (2.4). The following set (described above) in $F_{3}=F(\left\{x,y,z\right\})$ for $n=1$
$A=\{x,\ xz,\ y^{-1},\ y^{-1}x^{-1}y^{-1},\ y^{-1}z\},$
$A^{-1}=\{x^{-1},\ z^{-1}x^{-1},\ y,\ yxy,\ z^{-1}y\},$
has
$AA^{-1}~=~\{$ $e,\ xz^{-1}x^{-1},\ xy,\ xyxy,\ xz^{-1}y,\ xzx^{-1},\ xzy,\ xzyxy,\ y^{-1}x^{-1},\ y^{-1}z^{-1}x^{-1},\$
$y^{-1}z^{-1}y,\ y^{-1}x^{-1}y^{-1}x^{-1},\ y^{-1}x^{-1}y^{-1}z^{-1}x^{-1},\ y^{-1}x^{-1}y^{-1}z^{-1}y,$
$y^{-1}zx^{-1},\ y^{-1}zy,\ y^{-1}zyxy\},$
$A^{-1}A~=~\{$ $e,\ z,\ x^{-1}y^{-1},\ x^{-1}y^{-1}x^{-1}y^{-1},x^{-1}y^{-1}z,\ z^{-1},\ z^{-1}x^{-1}y^{-1},z^{-1}x^{-1}y^{-1}x^{-1}y^{-1},\$
$z^{-1}x^{-1}y^{-1}z,\ yx,\ yxz,\ yxyx,\ yxyxz,\ z^{-1}yx,\ z^{-1}yxz\}.$
Hence,
$|AA^{-1}|-|A^{-1}A|~=~17-15~=~2.$
More generally for $n\geq 1$, $A_{n}$ is constructed as a subset of $F_{3n}=F(\left\{x_{1},y_{1},z_{1},\dots,x_{n},y_{n},z_{n}\right\})$ as follows. Let
$A_{n}~\coloneqq~\bigcup_{i=1}^{n}\left\{x_{i},y_{i}^{-1},y_{i}^{-1}x_{i}^{-1}y_{i}^{-1},x_{i}z_{i},y_{i}^{-1}z_{i}\right\}.$
We claim
$|A_{n}A_{n}^{-1}|-|A_{n}^{-1}A_{n}|~=~2n.$
Define $A_{n}^{(i)}=\left\{x_{i},y_{i}^{-1},y_{i}^{-1}x_{i}^{-1}y_{i}^{-1},x_{i}z_{i},y_{i}^{-1}z_{i}\right\}$ for $1\leq i\leq n$. We have
$A_{n}A_{n}^{-1}$ $~=~\bigsqcup_{i=1}^{n}\bigsqcup_{j=1}^{n}A_{n}^{(i)}\left(A_{n}^{(j)}\right)^{-1},$
$A_{n}^{-1}A_{n}$ $~=~\bigsqcup_{i=1}^{n}\bigsqcup_{j=1}^{n}\left(A_{n}^{(i)}\right)^{-1}A_{n}^{(j)}.$
This implies [^2] For $i\neq j$, the sets $A_{n}^{(i)}$ and $A_{n}^{(j)}$ are supported on disjoint sets of generators in $F_{3k}$, meaning any product of the form $ab^{-1}$ or $a^{-1}b$ where $a\in A_{n}^{(i)}$ and $b\in A_{n}^{(j)}$ produces a word involving letters from distinct alphabets. Thus all such products are distinct. However, if $i=j$, the words produced come from the same alphabet, which doesn’t generate distinct products.
$|A_{n}A_{n}^{-1}|$ $~=~\sum_{i=1}^{n}\sum_{j=1}^{n}\left\lvert A_{n}^{(i)}\left(A_{n}^{(j)}\right)^{-1}\right\rvert$
$~=~\sum_{i=1}^{n}\left\lvert A_{n}^{(i)}\left(A_{n}^{(i)}\right)^{-1}\right\rvert+\sum_{i=1}^{n}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\left\lvert A_{n}^{(i)}\left(A_{n}^{(j)}\right)^{-1}\right\rvert$
$~=~n\cdot|AA^{-1}|+n\cdot(n-1)\cdot|A|^{2},$
and
$|A_{n}^{-1}A_{n}|$ $~=~\sum_{i=1}^{n}\sum_{j=1}^{n}\left\lvert\left(A_{n}^{(i)}\right)^{-1}A_{n}^{(j)}\right\rvert$
$~=~\sum_{i=1}^{n}\left\lvert\left(A_{n}^{(i)}\right)^{-1}A_{n}^{(i)}\right\rvert+\sum_{i=1}^{n}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\left\lvert\left(A_{n}^{(i)}\right)^{-1}A_{n}^{(j)}\right\rvert$
$~=~n\cdot|A^{-1}A|+n\cdot(n-1)\cdot|A|^{2}.$
Therefore,
$|A_{n}A_{n}^{-1}|-|A_{n}^{-1}A_{n}|~=~n\cdot(|AA^{-1}|-|A^{-1}A|)~=~2n.\qed$
Theorem 1.1 and Theorem 1.2 establish that every even integer arises in the difference between the left and right quotient set. Naturally, we may now ask how large such differences can be in terms of the cardinality of $A$. Define
$M_{n}~\coloneqq~\sup_{\begin{subarray}{c}A\subseteq F_{2}\\ |A|=n\end{subarray}}\left||AA^{-1}|-|A^{-1}A|\right|.$
Proposition 2.5.
We have $M_{n}=\Theta(n^{2})$.
Proof.
We establish both an upper and lower bound. For the lower bound, observe that for any finite subset $A\subseteq F_{2}$ with $|A|=n$ it follows that $|AA^{-1}|\leq n^{2}$ and $|A^{-1}A|\leq n^{2}$. This implies $||AA^{-1}|-|A^{-1}A||\leq n^{2}$ thus $M_{n}\leq n^{2}$ so $M_{n}=O(n^{2})$. For the lower bound, let
$A_{k}$ $~\coloneqq~\{x^{i}:1\leq i\leq k\},$
$B_{k}$ $~\coloneqq~\{x^{i}y:1\leq i\leq k\},$
$C_{k}$ $~\coloneqq~A_{k}\cup B_{k}.$
Note that $|C_{k}|=2k$ and thus let $n=2k$ . Next, we will compute the size of the right quotient set:
$C_{k}C_{k^{-1}}$ $~=~(A_{k}\cup B_{k})(B_{k}^{-1}\cup A_{k}^{-1})$
$~=~(A_{k}B_{k}^{-1})\cup(B_{k}B_{k}^{-1})\cup(A_{k}A_{k}^{-1})\cup(B_{k}A_{k}^{-1}).$
As $A_{k}A_{k}^{-1}=\{x^{i-j}\}=B_{k}B_{k}^{-1}$, this accounts for 2k12k-1 elements. As $A_{k}B_{k}^{-1}=\{x^{i}y^{-1}x^{-j}\}$, this adds $k^{2}$ elements, and lastly for $B_{k}A_{k}^{-1}=\{x^{i}yx^{-j}\}$, there are also $k^{2}$ elements. Thus,
$|C_{k}C_{k}^{-1}|~=~2k^{2}+2k-1.$
For the left quotient set, we obtain
$C_{k}^{-1}C_{k}$ $~=~(B_{k}^{-1}\cup A_{k}^{-1})(A_{k}\cup B_{k})$
$~=~(B_{k}^{-1}A_{k})\cup(B_{k}^{-1}B_{k})\cup(A_{k}^{-1}A_{k})\cup(A_{k}^{-1}B_{k}).$
As each of these four terms contribute 2k12k-1 elements,
$|C_{k}^{-1}C_{k}|~=~4(2k-1)=8k-4.$
Then, taking the difference between the left and right quotient set yields
$||C_{k}C_{k}^{-1}|-|C_{k}^{-1}C_{k}||$ $~=~(2k^{2}+2k-1)-(8k-4)$
$~=~2k^{2}-6k+1.$
Substituting $k=n/2$ gives $M_{n}\geq\frac{1}{2}n^{2}-3n+1=\Omega(n^{2})$. Thus, $M_{n}=\Theta(n^{2}).$
Remark 2.6.
The following remark is mentioned in [Tao08]: if $H\subseteq G$ is a finite, nonempty subset of some group $G$ and $g$ is an element of $G$ not in the normalizer of $H$, then the set $A\coloneqq Hg\cup H$ has $AA^{-1}$ about the same size as $H$, but $A^{-1}A$ can be possibly large. Our example in $F_{2}$ is this construction with $H=\left\{x,x^{2},\dots,x^{k}\right\}$ and $g=y$.
2.3. Cardinality of AA
The set $A$ given in (2.1) is extremal in the sense that it is the smallest such set such that $|AA^{-1}|-|A^{-1}A|$ in $F_{2}$. In fact, this holds for all groups with no elements of order $2$, which we prove in Theorem 1.3.
Lemma 2.7.
Let $G$ be a group with $A\subset G$ and $|A|=n$. Then $D_{A}$ contains no connected component (other then the diagonal) with more than $n$ elements.
Proof.
Suppose, for contradiction, $C$ was a connected component with more than $n$ elements. By the pigeonhole principle, two vertices in $C$ have the same first coordinate. Since $C$ is a clique by 2.1 (4), this contradicts 2.1 (2b). ∎
Lemma 2.8.
Let $G$ be a group with $A\subset G$ and $|A|=4$. If $G$ does not have an element of order 2 and the largest possible clique in $D_{A}$ is $K_{4}$, the number of connected components in $D_{A}$ is equal to the number of connected components in $D_{A^{-1}}$.
Proof.
Consider a triangle $\triangle=(\alpha_{1},\beta_{1})(\alpha_{2},\beta_{2})(\alpha_{3},\beta_{3})$. By 2.1 (2b), we know that the same number can appear at most twice in the coordinates $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$, $\alpha_{3}$, and $\beta_{3}$. Moreover the same number appears at most once for an $\alpha_{i}$ and at most once for a $\beta_{j}$, where $1\leq i,j\leq 3$. Therefore, there are two cases to consider.
Case 1: There are 4 distinct elements in $\triangle$. Because of the condition that there is no element of order 2 in $G$, there cannot be an edge between $(i,j)$ and $(j,i).$ Therefore, up to relabeling, $\triangle=(1,2)(2,3)(3,4)$. Then, $D_{A}$ and $D_{A^{-1}}$ look as follows.
$\phi$
Therefore, in this case, the number of connected components stays the same. Additionally there may be a triangle $\triangle^{\prime}=(3,1)(4,2)(1,4)$ and its reflection under $T$. However, we can check that the image of this graph under $\phi$ is itself as shown below.
$\phi$
Case 2: There are 3 distinct elements in $\triangle$. The only possibility in this case is that up to relabelling, $\triangle=(1,2)(2,3)(3,1).$ Notice that this triangle cannot be part of a $K_{4}$ because $(4,4)$ is not in the same connected component as $\triangle$ and every other coordinates will cause contradiction to Lemma 2.1 (2b). Then, $D_{A}$ and $D_{A^{-1}}$ are as follows.
$\phi$
In the second case, we see it is impossible to place a second triangle without two vertices being on the same axis, violating 2.1 (2b). ∎
Proof of Theorem 1.3.
If $|AA^{-1}|\neq|A^{-1}A|$, then the number of connected components in $D_{A}$ is different from $D_{A^{-1}}$. We claim at least one of $D_{A}$ and $D_{A}^{-1}$ contains a clique of size 3 or greater off the main diagonal (a non-diagonal clique). Indeed, since $\phi$ is a bijection between the non-diagonal and self-loop edges, the number of connected components in $D_{A}$ would be the same as $D_{A^{-1}}$ if there were no non-diagonal connected components of size 3 or greater.
Assume $D_{A}$ has a non-diagonal clique of size 3 or greater. This is impossible if $|A|\in\left\{1,2\right\}$ by 2.7. If $|A|=3$, then the only possible triangles are $(1,2)(3,1)(2,3)$ and $(2,1)(1,3)(3,2)$. Since any $D_{A}$ is invariant under transposition by $T$, if $D_{A}$ has a triangle, then it has both. Any other edge would either (1) connect the triangles to the diagonal, which is forbidden, or (2) connect the triangle to each other, which creates a component with more $3$ vertices, contradicting 2.7. Thus, with $|A|=3$, $D_{A}$ and $D_{A^{-1}}$ must have the same component counts. So we have shown for any group $G$ if $|A|\leq 3$ then $|AA^{-1}|=|A^{-1}A|$.
Notice how the above proof does not use 2.1 (2c), so it holds for any group, yielding the first part of the theorem.
Next, suppose that $G$ has no elements of order $2$. Then note that the size of any clique in $D_{A}$ is at most 4 by 2.7. We claim $D_{A}$ cannot have a $K_{4}$ other than the diagonal. Suppose $D_{A}$ did have such a $K_{4}$. Let the vertices of this $K_{4}$ be $(1,j_{1})$, $(2,j_{2})$, $(3,j_{3})$, $(4,j_{4})$. Up to relabeling, we may assume $j_{1}=2$. Further, we may assume up to relabeling that $j_{2}=3$. This forces $j_{4}\in\left\{1,4\right\}$, which implies it equals $1$. Therefore, $j_{3}=4$. This clique is impossible if $G$ has no elements of order $2$. Indeed,
$a_{1}a_{3}^{-1}~=~(a_{1}a_{2}^{-1})(a_{2}a_{3}^{-1})~=~(a_{3}a_{4}^{-1})(a_{4}a_{1}^{-1})~=~a_{3}a_{1}^{-1}.$
Therefore, $D_{A}$’s components are $K_{3}$’s and $K_{2}$’s. 2.8 implies that $\phi$ induces a bijection on the $K_{3}$’s, and hence on the $K_{2}$’s as well, by an edge counting argument combined with the fact that there are no $K_{4}$’s. Hence, we have shown if $G$ is a group with no elements of order 2 and $|A|\leq 4$, then $|AA^{-1}|=|A^{-1}A|$. ∎
Example 2.2.
If $G$ has elements of order $2$, then we may no longer use 2.1 (2c) in the previous statements. Indeed, we check with Sage [The25] that the set
[…]
which can be viewed as a subset of the quasidihedral group of order $16$, has $|AA^{-1}|=10$, but $|A^{-1}A|=7$. This simultaneously shows we can have an odd difference in the left and right quotient sets and also that $|A|=4$ can achieve a nonzero difference.
3. Future Work
In a finite subset $S$ of a group $G$, one may ask how many finite subsets $A\subseteq S$ have the left quotient set larger than the right quotient set. If $S$ is symmetric (i.e., $w\in S$ if and only if $w^{-1}\in S$), this is equal to the number of finite subsets with right quotient set larger than left quotient set. Indeed, if $|A^{-1}A|>|AA^{-1}|$, then we can replace $A$ with $B\coloneqq A^{-1}$ to get $|B^{-1}B|<|BB^{-1}|$, and vice-versa. In particular, this implies that the quantity
$\mathbb{E}[|A^{-1}A|-|AA^{-1}|]~=~0,$
where $A\subseteq S$ is chosen by including elements at random with probability $1/2$.
On the other hand, the second moment (i.e., the variance)
$\mathrm{Var}(|A^{-1}A|-|AA^{-1}|)~=~\mathbb{E}\left[(|A^{-1}A|-|AA^{-1}|)^{2}\right]$
is nonzero for groups where $|AA^{-1}|\neq|A^{-1}A|$ for some finite subset $A\subseteq S$. If $A$ is a subgroup then trivially one has the two cardinalities are the same, as in that case $A=A^{-1}$. Having to avoid such special cases is common; for example, there is non-Benford behavior if the support of certain random variables live in a coset of a group (see [MN08], in particular the discussion related to a result of Levy [Lév39]).
We can cast this problem specifically to the free group.
Question 3.1.
Let $B_{N}\subseteq F_{2}$ be the set of words of length $\leq N$. With respect to the uniform probability measure on the subsets of $B_{N}$, what is $\mathrm{Var}(|A^{-1}A|-|AA^{-1}|)$?
A technique for studying the moments of $|AA^{-1}|$ and related quantities is the graph theoretic framework in [LMO13]. This paper suggests that to study $\mathbb{P}(w_{1},\dots,w_{k}\notin AA^{-1})$, we should look at the graph with vertex set $S$ and edge relation $u\sim v$ if $uv^{-1}=w_{m}$ or $vu^{-1}=w_{m}$ for some $m=1,\dots,k$. It would be interesting to find a structure which encapsulates both the the above graphs and the graphs $D_{A}$ and $D_{A^{-1}}$ which we introduced.
There are still other unanswered questions about the values $|AA^{-1}|-|A^{-1}A|$ can take in various groups.
Question 3.2.
What are the necessary and sufficient conditions on a group $G$ so that there exists a finite subset $A\subseteq G$ such that $|A^{-1}A|\neq|AA^{-1}|$?
Of course, $G$ being non-abelian is a necessary condition for $|A^{-1}A|\neq|AA^{-1}|$. However, one can check that $|AA^{-1}|=|A^{-1}A|$ for all subsets $A$ of the symmetric group $S_{3}$. Hence, being non-abelian is not a sufficient condition. It is possible that $|AA^{-1}|=|A^{-1}A|$ is true for all $A\subseteq S_{3}$ only because $S_{3}$ is sufficiently small [^3] As $|S_{n}|=n!$ , the number of subsets is $2^{n!}$, which grows very rapidly; $2^{3!}=64$, $2^{4!}$ is approximately $1.7\cdot 10^{7}$, while $2^{5!}$ exceeds $10^{36}$; straightforward sampling cannot hope but to explore a very small percentage of possible $A$’s. [^4] A relevant fact here is that $|AA|<2|A|$ is a sufficient but not necessary condition for $|AA^{-1}|=|A^{-1}A|$ (see [Tao]). One consequence of this is that no set $A$ of size $|A|>|G|/2$ can have $|AA^{-1}|\neq|A^{-1}A|$. Since $|S_{3}|=6$, this means all subsets of $S_{3}$ where $|AA^{-1}|\neq|A^{-1}A|$ have size $|A|>|S_{3}|/2$ or $|A|\leq 3$, and by 1.3, this is enough to prove $|AA^{-1}|=|A^{-1}A|$ for all subsets $A\subseteq S_{3}$ only in terms of the sizes of subsets.. Therefore, to find a more robust example, we ask whether there are infinitely many non-abelian groups where $|A^{-1}A|\neq|AA^{-1}|$ is impossible.
Question 3.3.
Is there an infinite family of finite, non-isomorphic, non-abelian groups $\{G_{i}\}$ for which $|A^{-1}A|=|AA^{-1}|$ for all subsets $A$ of $G_{i}$?
An alternative way to characterize groups where $|AA^{-1}|\neq|A^{-1}A|$ is possible would be to find a finite collection of difference graphs which serve as minimal counterexamples to the statement $|AA^{-1}|=|A^{-1}A|$ for all $A\subseteq G$. We conjecture that such a collection exists. That is to say, there should be a finite collection of graphs $\{H_{1},\dots,H_{k}\}$ such that, if $|AA^{-1}|\neq|A^{-1}A|$ for some $A\subseteq G$, then there is some $B\subseteq G$ such that $D_{B}\cong H_{i}$ for some $i=1,\dots,k$. A counterexample to our conjecture would be an infinite sequence of groups $G_{\ell}$ such that for any $N>0$, there is some group $G_{m}$ in the collection for which $|AA^{-1}|=|A^{-1}A|$ for all sets $A$ of size <N<N, but $|AA^{-1}|\neq|A^{-1}A|$ for some $A$ of size $\geq N$.